OpenStudy (anonymous):

If cos^2(t) < 3/4 then sin^2(t) is a. =>0 but < 1/4 b. >1/4 but <= 1 c. >1 d. non e of the above

OpenStudy (jdoe0001):

$$ cos^2(t) \lt \cfrac{3}{4} \implies cos(t) \lt \sqrt{\cfrac{3}{4}} \implies cos(t) \lt \cfrac{\sqrt{3}}{2} \implies \cfrac{adj}{hyp}\\ b = \sqrt{hyp^2-adj^2} \implies \sqrt{4-3} = 1\\ -\cfrac{1}{2} \lt sin(t) \lt \cfrac{1}{2} \implies -\cfrac{1}{2} \lt sin^2(t) \lt \cfrac{1}{4} $$